# https://xiaojiu-z.github.io/Stable-Makeup.github.io/

### v1

@@ 3.2 Model Training and Inference
Specifically, during training, $\mathit{I_{s}}$ and $\mathit{I_{c}}$ obtained from the target image are fed into the content encoder and structural encoder, respectively, while the augmented $\mathit{I_{t}}$ is used as the input for the makeup encoder. This implies that the facial structures of the final generated target image are determined by the structural encoder, while the content is determined by the content encoder. The pre-trained U-Net is used to predict and remove the noise that needs to be eliminated for reconstructing the target image. By decoupling content and structure during training, we can control and adjust the content and structure separately, improving the accuracy and consistency of the generated images. During inference, $\mathit{I_{s}}$ and $\mathit{I_{c}}$ obtained from the source image are input to content encoder and structural encoder respectively.
3.2.3 Augmentation.
In our training pipeline, we have used a variety of augmentations, which are crucial for adapting to real-world scenarios and achieving successful makeup transfer. These augmentations can be divided into two categories: makeup augmentations and source augmentations. Makeup augmentations involves structural augmentations such as face warping and affine transformations applied to the input of the makeup encoder. The purpose of these augmentations is to eliminate Content information from the reference makeup image and to disrupt the pixel alignment between the reference makeup image and the source image. Source augmentations are designed to make our model adaptable to varying facial sizes and poses in the source images in real-world scenarios. These augmentations include synchronized affine transformations applied to both the source image and the target image.
3.2.4 Loss function.
@@
$L(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\mathbf{c_{i}},\mathbf{c_{e}},\mathbf{c_{m}}\right)\right\|_{2}^{2}\right],$ (1)
@@
4 Experiments
4.1 Experimental Setup
We employed Stable Diffusion V1-5 as the pre-trained diffusion model and trained it on the Makeup-real Dataset, which consists of 20k makeup-non-makeup pairs and each pair with the same identity. We focus on real-world conditions and utilize extremely aggressive augmentation, as described in Section 3.2. While training, the parameters of the D-P makeup encoder, makeup cross-attention layers, content encoder and structural encoder are updated, while the pre-trained text-to-image model remains frozen. The model underwent 100,000 iterations of training on 8 H800 GPUs, with a batch size of 16 per GPU and a learning rate of 5e-5. Inference was performed using DDIM as the sampler, with a step size of 30 and a guidance scale set to 1.5.
4.2 Evaluation Benchmarks
  • CPM-real[26] dataset: It is very diverse in terms of makeup styles, containing both color and pattern makeups. The degree of makeup can vary from light to heavy, from color-oriented to pattern-driven.
  • Makeup-wild[17] dataset: It contains facial images with various poses and expressions as well as complex backgrounds to test methods in the real-world environment.
4.3 Evaluation Metrics
To provide a thorough and objective assessment of the performance of each algorithm in different aspects of makeup transfer, we employed the following metrics: CLIP-I[29] and DINO-I[4] were utilized to assess makeup transfer performance. After makeup transfer, the content and structure information of the input image should not be changed. So we use SSIM [45] and L2-M to evaluate the degree of content and structural maintenance.
4.4 Comparison Methods
We performed comprehensive comparisons with the most representative makeup transfer algorithms currently available, including BeautyGAN[21], CPM[26], PSGAN[17], EleGANt[51], SCGAN[7], SPMT[59] and SSAT[40], and conduct comparisons on M-Wild dataset and CPM-real dataset respectively.
4.5 Experiment Results
4.5.1 Qualitative Comparison.
Figure 4: Qualitative comparison of different methods. Our results outperform other methods in terms of makeup detail transfer, ranging from light makeup to personalized heavy makeup.
Fig. 4 provides a visual comparison of our proposed Stable-Makeup method and other makeup transfer methods. To further evaluate the performance of these methods on various makeup styles, we used some popular makeup images from social media as reference makeup in addition to the test set. Overall, most methods maintained good consistency in content and structure, with only CPM introducing some artifacts during makeup transfer. In terms of makeup transfer capability, our method outperformed other methods by a significant margin. Other methods were only effective in transferring simple lip and eye shadow colors, and were only effective for light makeup. In contrast, our method not only performed well for light makeup, but also effectively aligned and transferred heavy makeup to the corresponding positions in the source image, while preserving makeup details and corresponding color distributions on the facial regions of the makeup image.
4.5.2 Quantitative Comparison.
Table 1: Quantitative comparison of different methods. Metrics that are bold and underlined represent methods that rank 1st and 2nd, respectively.
@@ 4.5 Experiment Results
BeautyGAN 0.818 0.927 0.878 48.775 0.775 0.913 0.858 53.323
CPM 0.815 0.923 0.862 108.374 0.793 0.921 0.837 114.288
PSGAN 0.739 0.929 0.850 48.816 0.693 0.915 0.839 50.137
EleGANt 0.827 0.928 0.891 46.939 0.787 0.917 0.872 48.995
SCGAN 0.816 0.927 0.878 47.595 0.776 0.915 0.851 50.062
SPMT 0.834 0.930 0.911 31.671 0.778 0.918 0.905 32.323
SSAT 0.823 0.931 0.901 34.512 0.773 0.910 0.881 35.205
Ours 0.860 0.935 0.906 31.483 0.825 0.928 0.883 32.255
@@
To test the ability of methods in the real-world environment, We use the non-makeup 334 images of the Makeup-Wild(M-Wild) dataset as source images, and randomly selected 334 images from the makeup images in the M-Wild dataset and the CPM-real dataset, respectively, as our reference makeup images.
@@
As shown in Table 1, It is evident that Stable-Makeup outperforms all other makeup transfer methods in terms of the DINO-I and CLIP-I metrics, indicating that our method has superior makeup transfer capabilities compared to other methods. Additionally, the SSIM and L2-M metrics achieved good scores when compared with other methods, indicating that our method is able to maintain good content and structure information in the transferred images. It should be noted that SSIM metrics may not accurately reflect the makeup transfer performance, as they may produce good scores when the transferred result is identical to the source image, indicating that the transfer algorithm did not work. In line with previous studies[26, 47, 59], we also listed it as a reference in our evaluation.
4.5.3 User Study.
Table 2: User study. Metrics that are bold represent method that rank 1st. Our method achieves the highest average score.
Method BeautyGAN CPM PSGAN EleGANt SCGAN SPMT SSAT Ours
Average Score 3.01 3.13 2.95 3.62 2.90 3.35 2.85 4.58
We conducted a user study to compare our method with other methods perceptually. We randomly select 10 source images and 10 reference images from both the M-Wild dataset and CPM-real dataset. For each evaluation, each user will see one source image with a reference makeup image, and one image generated by each method. 60 evaluators were asked to provide a comprehensive score for each generated image based on its makeup consistency, content and structure consistency, using a scale of 1 (worst) to 5 (best). The results of average scores in Table 2 indicate that our method outperforms the comparison methods. Our method achieves the highest average score, which clearly indicates that our method outperforms the comparison methods.
4.5.4 Ablation Study.
Figure 5: Qualitative ablation results. Figure (a) explores the classifier-free guidance scale. Figure (b) presents the ablation study of different training and inference settings.
@@
W.O. structure input 0.776 0.923 0.772 103.029
W.O. source aug 0.816 0.921 0.874 153.463
W.O decoupling training 0.806 0.927 0.791 32.856
W.O multi-layer 0.803 0.911 0.880 34.861
Ours full 0.825 0.928 0.883 32.255
Table 3: Quantitative ablation results on CPM-real dataset.
Firstly, we explored the effect of different parameters for the classifier-free guidance scale (GS) under makeup conditions in figure 5 (a). Next, in figure 5 (b), we conducted ablation experiments by systematically removing each module using the same experimental setup. As shown in figure 5 (a), it was found that a larger GS led to a greater degree of makeup transfer, while our method maintained good content and structural consistency of the source image under different guidance parameters. In figure 5 (b), it demonstrated that without the structural encoder input during inference (first column), the inconsistency in facial structure was observed, such as changes in the shape of eyebrows and mouth in the first column of the generated images. Without makeup augmentation during training (second column), the model does not possess the capability to transfer makeup, but rather, it solely transfers the facial region of the makeup image to the target image. Removing source augmentation during training (third column), the alignment and transferability of the model were weakened, and the background color was changed. When removing structural encoder during training (fourth column, W.O. decoupling training), the facial structure also could not be well preserved. Without the multi-layer strategy for the makeup encoder (fifth column), the transfer of makeup details was incomplete. When our method was fully implemented, it achieved a complete transfer of makeup details while maintaining good consistency of the content and structure of the source image.
@@
Quantitative ablation results are shown in Table 3. We performed quantitative ablation experiments using the same experimental setup on CPM-real dataset as Sec. 4.5. We observed that our method achieved optimal performance when all modules were retained, while removing the structural encoder input during inference (first row) or training (third row) led to a significant decrease in content and structure consistency. Removing the multi-layer strategy during training resulted in a significant decrease in makeup transfer ability, and the absence of source enhancement led to a significant decrease in background consistency. We exclude the setting without makeup augmentations, as shown in Figure 5 (b) second column, where the model can only perform face swapping and lacks the ability to perform makeup transfer. This could significantly affect the calculation of our makeup transfer metrics.
Figure 6: The visual results of Stable-Makeup demonstrate its robustness and generalizability for various applications.
4.6 More Applications
As illustrated in the Fig. 6, our Stable-Makeup not only achieves unprecedented makeup transfer effects on real faces, but also enables various applications that were previously unattainable with traditional makeup transfer methods. For instance, our method can perform makeup transfer on cross-domain faces and guide text-to-image generation with reference makeup conditions to create creative artistic images. Our method can also perform makeup transfer on videos by utilizing effective multi-frame concatenation. Moreover, our approach supports other domain’s reference makeup, such as animal-inspired subjects, animated characters and so on. providing a broader range of creative makeup options for real human faces.
5 Conclusion
In this paper, we propose the Stable-Makeup, a diffusion-based framework aiming for real-world makeup transfer. This method has made a significant breakthrough in the field of makeup transfer by achieving effects that were previously unattainable. The Stable-Makeup consists of detail-preserving makeup encoder, makeup cross-attention layers, content and structural control modules. The detail-preserving makeup encoder captures fine details of makeup, while two encoder structures are employed for the content and structural control modules to encode content and structure of source image. The makeup cross-attention layers effectively aligns facial region between reference makeup and source image. Through the content and structure decoupling training, we are able to further maintain content and structure consistency. In addition, we proposed a automatic pipeline to create a variety of makeup pairing data for training. Our proposed method has surpassed all existing makeup transfer algorithms and has expanded the boundaries of the makeup transfer field. This breakthrough has significant implications for the practical application of makeup transfer technology in various fields, such as cosmetics, entertainment, and fashion.
Supplementary Materials
F Preliminaries
F.0.1 Diffusion models.
Diffusion Model (DM) [13, 39] belongs to the category of generative models that denoise from a Gaussian prior $\mathbf{x_{T}}$ to target data distribution $\mathbf{x_{0}}$ by means of an iterative denoising procedure. Latent Diffusion Model (LDM) [31] is proposed to model image representations in autoencoder’s latent space. LDM significantly speeds up the sampling process and facilitates text-to-image generation by incorporating additional text conditions. The LDM loss is:
$L_{LDM}(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\bm{\tau_{\theta}}(\mathbf{c})\right)\right\|_{2}^{2}\right],$ (2)
[…]
Stable Diffusion (SD) is a widely adopted text-to-image diffusion model based on LDM. Compared to LDM, SD is trained on a large LAION [36] dataset and replaces BERT with the pre-trained CLIP [29] text encoder.
F.0.2 ControlNet.
ControlNet [54] is one of the most popular control modules in current diffusion models. It receives inputs from various modalities as spatial control conditions, guiding the diffusion model to generate desired images, thereby achieving controllable generation. ControlNet copies an identical U-Net structure as trainable parameters and locks the original U-Net parameters, The trainable copy is connected to the locked model with zero convolution layers, denoted $\mathcal{Z}(\cdot;\cdot)$. The complete ControlNet can represented as:
$\boldsymbol{y}_{\mathrm{c}}=\mathcal{F}(\boldsymbol{x};\Theta)+\mathcal{Z}\left(\mathcal{F}\left(\boldsymbol{x}+\mathcal{Z}\left(\boldsymbol{c};\Theta_{\mathrm{z}1}\right);\Theta_{\mathrm{c}}\right);\Theta_{\mathrm{z}2}\right)$ (3)
Where $\Theta_{\mathrm{z}1}$ and $\Theta_{\mathrm{z}2}$ represent two different zero conv layers’ parameters respectively. $\mathcal{F}$ represent diffusion U-Net, $\boldsymbol{x}$ represent image latent, $\Theta$ is the frozen weight of the U-Net and $\Theta_{\mathrm{c}}$ is the trainable copy weight of the U-Net.
F.0.3 CLIP.
CLIP [29] consists of two integral components: an image encoder, represented as $F(x)$, and a text encoder, represented as $G(t)$. The image encoder, $F(x)$, transforms an image $x$ with dimensions $\mathbb{R}^{3\times H\times W}$ (height $H$ and width $W$) into a $d$-dimensional image feature $f_{x}$ with dimensions $\mathbb{R}^{N\times d}$, where $N$ is the number of divided patches. On the other hand, the text encoder, $G(t)$, creates a $d$-dimensional text representation gt with dimensions $\mathbb{R}^{M\times d}$ from natural language text $t$, where $M$ is the number of text prompts. After training on a contrastive loss function, CLIP can be applied directly for zero-shot image recognition without the need for fine-tuning the entire model.
G Further Exploration of Data Construction

### v2

@@ 3.2. Model Training and Inference
Specifically, during training, $\mathit{I_{s}}$ and $\mathit{I_{c}}$ obtained from the target image are fed into the content encoder and structural encoder, respectively, while the augmented $\mathit{I_{t}}$ is used as the input for the makeup encoder. This implies that the facial structures of the final generated target image are determined by the structural encoder, while the content is determined by the content encoder. The pre-trained U-Net is used to predict and remove the noise that needs to be eliminated for reconstructing the target image. By decoupling content and structure during training, we can control and adjust the content and structure separately, improving the accuracy and consistency of the generated images. During inference, $\mathit{I_{s}}$ and $\mathit{I_{c}}$ obtained from the source image are input to the content encoder and structural encoder respectively.
3.2.3. Augmentation.
In our training pipeline, we have used a variety of augmentations, which are crucial for adapting to real-world scenarios and achieving successful makeup transfer. These augmentations can be divided into two categories: makeup augmentations and source augmentations. Makeup augmentations involve structural augmentations such as face warping and affine transformations applied to the input of the makeup encoder. The purpose of these augmentations is to eliminate Content information from the reference makeup image and to disrupt the pixel alignment between the reference makeup image and the source image. Source augmentations are designed to make our model adaptable to varying facial sizes and poses in the source images in real-world scenarios. These augmentations include synchronized affine transformations applied to both the source image and the target image.
3.2.4. Loss function.
@@
(1) $L(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\mathbf{c_{i}},\mathbf{c_{e}},\mathbf{c_{m}}\right)\right\|_{2}^{2}\right],$
@@
Figure 5. Qualitative ablation results. Figure (a) explores the classifier-free guidance scale. Figure (b) presents the ablation study of different training and inference settings.
4. Experiments
@@
BeautyGAN 0.742 0.926 0.884 48.408 0.718 0.915 0.866 49.168
CPM 0.754 0.920 0.873 94.092 0.757 0.920 0.846 95.172
PSGAN 0.748 0.928 0.862 45.742 0.724 0.916 0.829 46.091
EleGANt 0.759 0.928 0.878 44.463 0.731 0.917 0.838 45.797
SCGAN 0.751 0.928 0.882 44.296 0.726 0.916 0.858 46.284
SPMT 0.761 0.931 0.915 29.126 0.735 0.920 0.900 28.299
SSAT 0.758 0.930 0.914 32.578 0.719 0.915 0.896 33.604
Ours 0.768 0.933 0.910 29.087 0.760 0.923 0.868 27.976
Table 1. Quantitative comparison of different methods. Metrics that are red and blue represent methods that rank 1st and 2nd, respectively.
4.1. Experimental Setup
We employed Stable Diffusion V1-5 as the pre-trained diffusion model and trained it on the Makeup-real Dataset, which consists of 20k makeup-non-makeup pairs, and each pair with the same identity. We focus on real-world conditions and utilize extremely aggressive augmentation. While training, the parameters of the D-P makeup encoder, makeup cross-attention layers, content encoder, and structural encoder are updated, while the pre-trained text-to-image model remains frozen. The model underwent 100,000 iterations of training on 8 H800 GPUs, with a batch size of 16 per GPU and a learning rate of 5e-5. Inference was performed using DDIM as the sampler, with a step size of 30 and a guidance scale set to 1.5.
4.2. Evaluation Metrics
To provide a thorough and objective assessment of the performance of each algorithm in different aspects of makeup transfer, we employed the following metrics: CLIP-I(Radford et al., 2021) and DINO-I(Caron et al., 2021) were utilized to assess makeup transfer performance. After makeup transfer, the content and structure information of the input image should not be changed. So we use SSIM (Wang et al., 2004) and L2-M to evaluate the degree of content and structural maintenance.
4.3. Comparison Methods
We performed comprehensive comparisons with the most representative makeup transfer algorithms currently available, including BeautyGAN(Li et al., 2018), CPM(Nguyen et al., 2021), PSGAN(Jiang et al., 2020), EleGANt(Yang et al., 2022), SCGAN(Deng et al., 2021), SPMT(Zhu et al., 2022) and SSAT(Sun et al., 2023), and conduct comparisons on M-Wild dataset(Jiang et al., 2020) and CPM-real dataset(Nguyen et al., 2021) respectively.
4.4. Experiment Results
4.4.1. Qualitative Comparison.
Figure 6. The visual results of Stable-Makeup demonstrate its robustness and generalizability for various applications.
Fig. 4 provides a visual comparison of our proposed Stable-Makeup method and other makeup transfer methods. To further evaluate the performance of these methods on various makeup styles, we used some popular makeup images from social media as reference makeup in addition to the test set. Overall, most methods maintained good consistency in content and structure, with only CPM introducing some artifacts during makeup transfer. In terms of makeup transfer capability, our method outperformed other methods by a significant margin. Other methods were only effective in transferring simple lip and eye shadow colors and were only effective for light makeup. In contrast, our method not only performed well for light makeup but also effectively aligned and transferred heavy makeup to the corresponding positions in the source image, while preserving makeup details and corresponding color distributions on the facial regions of the makeup image.
4.4.2. Quantitative Comparison.
@@ 4.4. Experiment Results
To test the ability of methods in the real-world environment, We utilized all makeup images from M-wild Dataset and CPM-Real dataset as reference images, while randomly selecting non-makeup images from the MT-dataset and MWild-Dataset as source images. We utilized a total of 7017 makeup reference images, paired with an equal number of source images.
@@
As shown in Table 1, It is evident that Stable-Makeup outperforms all other makeup transfer methods in terms of the DINO-I and CLIP-I metrics, indicating that our method has superior makeup transfer capabilities compared to other methods. Additionally, the SSIM and L2-M metrics achieved good scores when compared with other methods, indicating that our method is able to maintain good content and structure information in the transferred images. It should be noted that SSIM metrics may not accurately reflect the makeup transfer performance, as they may produce good scores when the transferred result is identical to the source image, indicating that the transfer algorithm did not work. In line with previous studies(Nguyen et al., 2021; Xiang et al., 2022; Zhu et al., 2022), we also listed it as a reference in our evaluation.
4.4.3. Ablation Study.
Firstly, we explored the effect of different parameters for the classifier-free guidance scale (GS) under makeup conditions in figure 5 (a). Next, in figure 5 (b), we conducted ablation experiments by systematically removing each module using the same experimental setup. As shown in figure 5 (a), it was found that a larger GS led to a greater degree of makeup transfer, while our method maintained good content and structural consistency of the source image under different guidance parameters. In figure 5 (b), it demonstrated that without the structural encoder input during inference (first column), the inconsistency in facial structure was observed, such as changes in the shape of eyebrows and mouth in the first column of the generated images. Without makeup augmentation during training (second column), the model does not possess the capability to transfer makeup, but rather, it solely transfers the facial region of the makeup image to the target image. Removing source augmentation during training (third column), the alignment and transferability of the model were weakened, and the background color was changed. When removing the structural encoder during training (fourth column, W.O. decoupling training), the facial structure also could not be well preserved. Without the multi-layer strategy for the makeup encoder (fifth column), the transfer of makeup details was incomplete. When our method was fully implemented, it achieved a complete transfer of makeup details while maintaining good consistency of the content and structure of the source image. Furthermore, we provide additional quantitative ablation results in supplementary.
4.5. More Applications
As illustrated in Fig. 6, our Stable-Makeup not only achieves unprecedented makeup transfer effects on real faces but also enables various applications that were previously unattainable with traditional makeup transfer methods. For instance, our method can perform makeup transfer on cross-domain faces and guide text-to-image generation with reference makeup conditions to create creative artistic images. Our method can also perform makeup transfer on videos by utilizing effective multi-frame concatenation. Moreover, our approach supports other domain’s reference makeup, such as animal-inspired subjects, animated characters and so on. providing a broader range of creative makeup options for real human faces.
5. Conclusion
In this paper, we propose the Stable-Makeup, a novel diffusion-based framework aiming for real-world makeup transfer. This method has made a significant breakthrough in the field of makeup transfer by achieving effects that were previously unattainable. The Stable-Makeup consists of three core modules. The detail-preserving makeup encoder captures intricate makeup details, integrating with two modules for content and structural control of the source image. Makeup cross-attention layers align facial regions between reference and source images, while content and structure decoupling training enhances consistency. Additionally, we propose an automated pipeline for generating diverse makeup pairing data for training. Our proposed method has surpassed all existing makeup transfer algorithms and has expanded the boundaries of the makeup transfer field. This breakthrough has significant implications for the practical application of makeup transfer technology in various fields.
Supplementary Materials
Appendix A Preliminaries
A.0.1. Diffusion models.
Diffusion Model (DM) (Ho et al., 2020; Song et al., 2020) belongs to the category of generative models that denoise from a Gaussian prior $\mathbf{x_{T}}$ to target data distribution $\mathbf{x_{0}}$ by means of an iterative denoising procedure. Latent Diffusion Model (LDM) (Rombach et al., 2022) is proposed to model image representations in the autoencoder’s latent space. LDM significantly speeds up the sampling process and facilitates text-to-image generation by incorporating additional text conditions. The LDM loss is:
(2) $L_{LDM}(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\bm{\tau_{\theta}}(\mathbf{c})\right)\right\|_{2}^{2}\right],$
[…]
Stable Diffusion (SD) is a widely adopted text-to-image diffusion model based on LDM. Compared to LDM, SD is trained on a large LAION (Schuhmann et al., 2022) dataset and replaces BERT with the pre-trained CLIP (Radford et al., 2021) text encoder.
A.0.2. ControlNet.
ControlNet (Zhang et al., 2023b) is one of the most popular control modules in current diffusion models. It receives inputs from various modalities as spatial control conditions, guiding the diffusion model to generate desired images, thereby achieving controllable generation. ControlNet copies an identical U-Net structure as trainable parameters and locks the original U-Net parameters, The trainable copy is connected to the locked model with zero convolution layers, denoted $\mathcal{Z}(\cdot;\cdot)$. The complete ControlNet can represented as:
(3) $\bm{y}_{\mathrm{c}}=\mathcal{F}(\bm{x};\Theta)+\mathcal{Z}\left(\mathcal{F}\left(\bm{x}+\mathcal{Z}\left(\bm{c};\Theta_{\mathrm{z}1}\right);\Theta_{\mathrm{c}}\right);\Theta_{\mathrm{z}2}\right)$
Where $\Theta_{\mathrm{z}1}$ and $\Theta_{\mathrm{z}2}$ represent two different zero conv layers’ parameters respectively. $\mathcal{F}$ represent diffusion U-Net, $\bm{x}$ represent image latent, $\Theta$ is the frozen weight of the U-Net and $\Theta_{\mathrm{c}}$ is the trainable copy weight of the U-Net.
A.0.3. CLIP.
CLIP (Radford et al., 2021) consists of two integral components: an image encoder, represented as $F(x)$, and a text encoder, represented as $G(t)$. The image encoder, $F(x)$, transforms an image $x$ with dimensions $\mathbb{R}^{3\times H\times W}$ (height $H$ and width $W$) into a $d$-dimensional image feature $f_{x}$ with dimensions $\mathbb{R}^{N\times d}$, where $N$ is the number of divided patches. On the other hand, the text encoder, $G(t)$, creates a $d$-dimensional text representation gt with dimensions $\mathbb{R}^{M\times d}$ from natural language text $t$, where $M$ is the number of text prompts. After training on a contrastive loss function, CLIP can be applied directly for zero-shot image recognition without the need for fine-tuning the entire model.
Appendix B Evaluation Benchmarks
  • CPM-real dataset: It is very diverse in terms of makeup styles, containing both color and pattern makeup. The degree of makeup can vary from light to heavy, from color-oriented to pattern-driven.
  • M-wild dataset: It contains facial images with various poses and expressions as well as complex backgrounds to test methods in a real-world environment.
@@ Appendix B Evaluation Benchmarks
W.O. structure input 0.716 0.913 0.759 97.746
W.O. source aug 0.752 0.911 0.858 96.385
W.O decoupling training 0.745 0.917 0.754 35.675
W.O multi-layer 0.742 0.907 0.861 29.076
Ours full 0.760 0.923 0.868 27.976
Table 2. Quantitative ablation results on CPM-real dataset.
Appendix C Quantitative Ablation
@@
Quantitative ablation results are shown in Table 2. We performed quantitative ablation experiments using the same experimental setup on the CPM-real dataset as the experiments section. We observed that our method achieved optimal performance when all modules were retained while removing the structural encoder input during inference (first row) or training (third row) led to a significant decrease in content and structure consistency. Removing the multi-layer strategy during training resulted in a significant decrease in makeup transfer ability, and the absence of source enhancement led to a significant decrease in background consistency. We exclude the setting without makeup augmentations, as shown in Figure 5 (b) second column in our paper, where the model can only perform face swapping and lacks the ability to perform makeup transfer. This could significantly affect the calculation of our makeup transfer metrics.
Appendix D Further Exploration of Data Construction